OpenStudy (anonymous):

What is the inverse of the function f(x) = 4x + 8?

OpenStudy (anonymous):

@Mehek14

Nnesha (nnesha):

f(x) is same as y so first replace f(x) with y \[\huge\rm \color{ReD}{y}=4\color{blue}{x}+8\] to find inverse `switch x and y` \[\huge\rm \color{blue}{x}=4\color{red}{y}+8\]now solve for y

OpenStudy (anonymous):

a. h(x) = 1/4x – 2 b. h(x) = 1/4x + 2 c. h(x) = 1/2x – 2 d. h(x) = 1/2x + 2 @Nnesha

Nnesha (nnesha):

ok. solve the equation for y ^^

OpenStudy (anonymous):

4 @Nnesha

Nnesha (nnesha):

:hmm there supposed to be x variable fist move 8 to the left side (subtract 8 both sides) and then move 4 to the left side by dividing 4 both sides :=)